Question 4: General Motors is thinking about t hree years Can you explain use the three nolons t whether
Answer #1

Answer:

Initial Investment =$365 millions

Yield=$420 millions

r=4%

n=3 years

So PV of Investment= 420/(1+4%)^3 - 365=$8.37 millions

FV of Investment=420-365*(1+4%)^3=$9.42 millions

let i is the IRR

420/(1+i)^3=365

i=4.79%